# Characterizing higher Auslander(-Gorenstein) Algebras

### v1

@@ Abstract.
Let $n$ be a positive integer. In this paper, by using torsion theoretic methods, we show that $n$-Auslander algebras can be characterized by the abelianness of the category of modules with projective dimension less than $n$ and a certain additional property, extending the classical Auslander-Tachikawa theorem. By Auslander-Iyama correspondence a categorical characterization of the class of Artin algebras having $n$-cluster tilting modules is obtained.
Since higher Auslander algebras are a special case of higher Auslander-Gorenstein algebras, the results are given in the general setting as extending previous results of Kong. Moreover, as an application of some results, we give categorical descriptions for the semisimplicity and selfinjectivity of an Artin algebra.
@@ Introduction
In 1971, Auslander [Au71] proved a remarkable classical result, called Auslander correspondence, about algebras of finite representation type which introduced a completely new insight to the representation theory of Artin algebras. He showed that there is a bijective correspondence between the set of Morita equivalence classes of Artin algebras $\Lambda$ of finite representation type and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an Auslander algebra; i.e. i.e. Artin algebras $\Gamma$ whose global dimensions (denoted by ${\rm{gldim}}(\Gamma)$) are at most two and whose dominant dimensions (denoted by ${\rm{domdim}}(\Gamma)$) are at least two. See [Au71, Chapter III, Section 4, Page 52] or [ARS, Theorem VI.5.7].
For an Artin algebra $\Gamma$, let ${\rm{Mod}}(\Gamma)$ be the category of all left $\Gamma$-modules and ${\rm{{mod}}}(\Gamma)$ be the category of all finitely generated left $\Gamma$-modules. If $\Gamma$ is an Auslander algebra, then the full subcategory $\mathrm{proj}(\Gamma)$ of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated projective $\Gamma$-modules is equivalent to a module category and so is itself an abelian category [Au71, §III, Section 4, Page 52]. This gave Tachikawa the motivation to study rings $\Gamma$ such that the full subcategory ${\mathrm{Proj}}(\Gamma)$ of ${\rm{Mod}}(\Gamma)$ consisting of all projective $\Gamma$-modules is abelian. He obtained some characterizations of such rings and proved that these algebras are precisely Auslander algebras [Ta74, Theorem 1]. For Artin algebras, by combining Auslander’s result and Tachikawa’s result, we have the following characterization of Auslander algebras.
@@
Hence, for Artin algebras, by Auslander correspondence and Auslander-Tachikawa theorem, there are bijections between the following four sets:
  • (i)
    Morita equivalence classes of Artin algebras $\Lambda$ such that $\Lambda$ is of finite representation type;
  • (ii)
    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an Auslander algebra;
  • (iii)
    Morita equivalence classes of Artin algebras $\Delta$ such that $\mathrm{proj}(\Delta)$ is an abelian category;
  • (iv)
    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{Proj}}(\Sigma)$ is an abelian category.
Moreover, an Artin algebra $\Lambda$ is of finite representation type if and only if ${\rm{{mod}}}(\Lambda)\simeq\mathrm{proj}(\Gamma)$, where $\Gamma$ is an Auslander algebra; see also [Iy04, Theorem 3.3].
In 2007, Iyama in a series of important papers [Iy07a, Iy07b] gave a generalization of Auslander correspondence, called higher Auslander correspondence or Auslander-Iyama correspondence. He showed that for every positive integer $n$, there is a bijective correspondence between the set of equivalence classes of $n$-cluster tilting subcategories with additive generators (Definition 5.5) of Artin algebras and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that ${\rm{gldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$, called $n$-Auslander algebras [Iy07b, Theorem 0.2].
A natural question is whether there exists a characterization for higher Auslander algebras analogous to Auslander-Tachikawa Theorem. The first goal of this paper is to answer this question and we have the following characterization for higher Auslander algebras; see Theorem 5.1.
@@
Here, for an Artin algebra $\Gamma$, we denote by $\Gamma^{{\rm{op}}}$ the opposite algebra of $\Gamma$, by ${\mathrm{Inj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all injective $\Gamma$-modules, by $\mathrm{inj}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated injective $\Gamma$-modules; for every non-negative integer $m$, by ${\mathrm{Proj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of projective dimension at most $m$, and by $\mathrm{proj}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of projective dimension at most $m$. For $m\geq 1$, we write
[…]
@@
[…]
@@
Note that only the abelianness property of the category of modules with projective dimension less than $n$ can not characterize $n$-Auslander algebras. For $n\geq 2$, the other conditions are necessary, although they hold automatically for $n=1$; see Remarks 5.2 and 5.3.
In 1993, Auslander and Solberg [AS93] established a Gorenstein analogue of Auslander correspondence by using the technique of relative homology developed in [AS93a, AS93b, AS93c]. They introduced $\tau$-selfinjective algebras, where $\tau$ is the Auslander-Reiten translation, as a replacement of representation-finite algebras. The corresponding analogue of Auslander algebras are characterized as Artin algebras whose injective dimensions are at most $2$ and whose dominant dimensions are at least $2$, that is, Artin algebras $\Gamma$ such that ${\rm{id}}({}_{\Gamma}\Gamma)\leq 2\leq\rm{domdim}(\Gamma)$. Kong named this class of algebras as quasi-Auslander algebras [Ko14]. In his work [Ko14] Kong also proved an extension of Auslander-Tachikawa’s theorem by characterizing quasi-Auslander algebras as Artin algebras over which the category of finitely generated Gorenstein projective modules is an abelian category.
Hence, for Artin algebras, by Auslander-Solberg correspondence ([AS93, Proposition 5.11]) and Kong’s theorem [Ko14, Theorem 2.1], there are bijections between the following four sets:
  • (i)
    Equivalence classes of suitable modules of Artin algebras;
  • (ii)
    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is a quasi-Auslander algebra;
  • (iii)
    Morita equivalence classes of Artin algebras $\Delta$ such that ${\mathrm{Gproj}}(\Delta)$ is an abelian category;
  • (iv)
    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{GProj}}(\Sigma)$ is an abelian category.
To obtain a higher analogue of Auslander-Solberg correspondence, Iyama and Solberg [IS18] introduced pre-cluster tilting subcategories in module categories of Artin algebras which are called $\tau_{n}$-selfinjective algebras as a higher version of $\tau$-selfjinjective algebras. They also gave the notion of $n$-minimal Auslander-Gorenstein algebras as a higher version of quasi-Auslander algebras, say Artin algebras $\Gamma$ with ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$. They showed that there exists a bijective correspondence between Morita equivalence classes of $n$-minimal Auslander-Gorenstein algebras and equivalence classes of finite $n$-precluster tilting subcategories of Artin algebras.
In view of the above Iyama-Solberg correspondence, we could establish a Gorenstein version of Theorem 2 which are also higher versions of Kong’s result [Ko14, Theorem 2.1].
@@
Here for every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$, by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
Hence, for Artin algebras, by Iyama-Solberg correspondence [IS18, Theorem 4.5] and our results (Theorem 2.6), there are bijections between the following sets:
  • (I)
    Equivalence classes of finite $n$-precluster tilting subcategories of Artin algebras;
  • (II)
    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra;
  • (III)
    Morita equivalence classes of Artin algebras $\Delta$ such that ${\mathrm{Gproj}}^{\leq n-1}(\Delta)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Delta)\cap\mathrm{inj}(\Delta)\subseteq\mathrm{proj}(\Delta)$, and ${{}^{\perp_{0}}\Delta}\subseteq{{}^{\perp_{n}}\Delta}$;
  • (IV)
    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{GProj}}^{\leq n-1}(\Sigma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Sigma)\cap{\mathrm{Inj}}(\Sigma)\subseteq{\mathrm{Proj}}(\Sigma)$, and ${{}^{\underline{\perp}_{0}}\Sigma}\subseteq{{}^{\underline{\perp}_{n}}\Sigma}$.
While the results of Section 2 mainly make use of torsion theoretic methods, in Section 4, we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS]. Especially, by Corollary 4.4, if $n$ is a positive integer and $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, then we have the following equivalences:
@@ 1.1. Conventions
Let $\Gamma$ be an Artin algebra defined over $R$. We denote by $\mathbf{D}$ the canonical duality ${\rm{Hom}}_{R}(-,\mathbf{E}$ $(R/J_{R})):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$, where $\mathbf{E}(R/J_{R})$ is the injective envelop of $R/J_{R}$ as $R$-module.
@@
For a module $M\in{\rm{{mod}}}(\Gamma)$ and $i\geq 0$, write $\mathrm{sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{add}(M))$, $\mathrm{fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{add}(M))$, $\mathrm{Sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{Add}(M))$, $\mathrm{Fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{Add}(M))$. In particular, $\mathrm{sub}(M):=\mathrm{sub}^{1}(M)$, $\mathrm{fac}(M)$ $:=\mathrm{fac}_{1}(M)$, $\mathrm{Sub}(M):=\mathrm{Sub}^{1}(M)$, $\mathrm{Fac}(M):=\mathrm{Fac}_{1}(M)$.
@@ 1.2. Dominant Dimension
The left dominant dimension of the ring $\Gamma$ is defined as the dominant dimension of the left regular module ${{}_{\Gamma}\Gamma}$. Note that the dominant dimension of a right module and the right dominant dimension of a ring are defined similarly. It is well-known that ${\rm{domdim}}({{}_{\Gamma}\Gamma})={\rm{domdim}}(\Gamma_{\Gamma})$ ; see [Mu68, Theorem 4]. So for the rest of the paper, we will denote both left and right dominant dimension of $\Gamma$, by ${\rm{domdim}}(\Gamma)$ and call it the dominant dimension of $\Gamma$.
@@
The inclusion ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$ can be proves exactly in the same way. ∎
@@ 1.3. Torsion Pairs
Let $\mathcal{A}$ be an abelian category. If $X,Y\in\mathcal{A}$, then we denote the set of morphisms from $X$ to $Y$ in $\mathcal{A}$ by $\mathcal{A}(X,Y)$ or ${\rm{Hom}}_{\mathcal{A}}(X,Y)$. All subcategories considered are assumed to be full and closed under isomorphisms and finite direct sums.
A torsion pair (or torsion theory) in the abelian category $\mathcal{A}$ is a pair $(\mathcal{X},\mathcal{Y})$ of subcategories such that $\mathcal{A}(X,Y)=0$ for all $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$, and these two classes are maximal for this property [Di66, Sections 1 and 2], that is,
@@
A subcategory of $\mathcal{A}$ is the torsion class (resp. the torsion-free class) of some torsion pair if and only if it is closed under quotients, direct sums and extensions (resp. subobjects, direct products and extensions, respectively) [Di66, Theorem 2.3].
If $(\mathcal{X},\mathcal{Y})$ is a torsion pair in $\mathcal{A}$, to each object $A$ in $\mathcal{A}$ is associated a unique short exact sequence (called the canonical sequence for $A$)
$0\longrightarrow xA\longrightarrow A\longrightarrow yA\longrightarrow 0$
such that $xA\in\mathcal{X}$ and $yA\in\mathcal{Y}$. In fact, $xA$ is the maximal subobject of $A$ contained in $\mathcal{X}$ [Di66, Proposition 2.4]. Also, the canonical sequence is functorial, i.e. it defines the functors $x:\mathcal{A}\longrightarrow\mathcal{X}$ given by $A\mapsto xA$ and $y:\mathcal{A}\longrightarrow\mathcal{Y}$ given by $A\mapsto yA$, see [BR07, Section 1.1]. Note that the composition of the functor $x:\mathcal{A}\longrightarrow\mathcal{X}$ by the inclusion functor $\mathcal{X}\longrightarrow\mathcal{A}$ is an idempotent radical, see [St75, Proposition VI.2.3].
A torsion pair $(\mathcal{X},\mathcal{Y})$ in $\mathcal{A}$ is called hereditary if $\mathcal{X}$ is also hereditary, i.e. $\mathcal{X}$ is closed under subobjects (which is equivalent to $\mathcal{Y}$ being closed under taking injective envelopes [Di66, Theorem 2.9]).
Let now $\Gamma$ be an Artin algebra. We would like to mention that in this case, a subcategory of ${\rm{{mod}}}(\Gamma)$ is the torsion class (resp. the torsion-free class) of some torsion pair if and only if it is closed under quotients and extensions (resp. submodules and extensions, respectively). For more details about torsion theories in ${\rm{{mod}}}(\Gamma)$; see [AIR, Subsection 1.1] or [Iy04, Subsection 1.2].
1.4. Functorially Finite Subcategories
Let $\Gamma$ be an Artin algebra. A subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Gamma)$ is said to be covariantly finite in ${\rm{{mod}}}(\Gamma)$, if all representable covariant functors on ${\rm{{mod}}}(\Gamma)$ are finitely generated when considered as functors on $\mathcal{X}$ by restriction. In other words, for each $M\in{\rm{{mod}}}(\Gamma)$, there exist $X\in\mathcal{X}$ and a morphism $\varphi:M\longrightarrow X$ so that the sequence
${\rm{Hom}}_{\Gamma}(X,-)\mid_{\mathcal{X}}\longrightarrow{\rm{Hom}}_{\Gamma}(M,-)\mid_{\mathcal{X}}\longrightarrow 0$
is exact. Here, the notation $\mid_{\mathcal{X}}$ means restriction to $\mathcal{X}$.
Dually, $\mathcal{X}$ is said to be contravariantly finite in ${\rm{{mod}}}(\Gamma)$ if the representable contravariant functors on ${\rm{{mod}}}(\Gamma)$ stay finitely generated when restricted to $\mathcal{X}$.
A subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Gamma)$ is said to be functorially finite if it is both contravariantly and covariantly finite in ${\rm{{mod}}}(\Gamma)$.
Note that if the pair $(\mathcal{X},\mathcal{Y})$ is a torsion pair in ${\rm{{mod}}}(\Gamma)$, then an easy investigation shows that $\mathcal{X}$ is a contravariantly finite and $\mathcal{Y}$ is covariantly finite in ${\rm{{mod}}}(\Gamma)$ [Sm84, Page 519].
@@ 1.4. Functorially Finite Subcategories
We recall some definitions and notations about (co)tilting modules and their relations to torsion theory; see [Ho82, As84, Sm84, AIR].
@@
Let $\Gamma$ be an Artin algebra. A $\Gamma$-module $T$ is said to be a tilting module, if the following three properties are satisfied.
@@
    There is an exact sequence $0\longrightarrow\Gamma\longrightarrow T^{\prime}\longrightarrow T^{\prime\prime}\longrightarrow 0$ with $T^{\prime}$ and $T^{\prime\prime}$ direct summands of direct sums of copies of $T$.
Dually, the $\Gamma$-module $T$ will be called a cotilting module if $\mathbf{D}(T)$ is a tilting $\Gamma^{\rm op}$-module.
@@
For the convenience of the reader, we recall the following relevant result on torsion pairs that we will use of it in next sections.
Theorem 1.4 ([Sm84, Theorem] and [AIR, Proposition 1.1]).
Let $\Gamma$ be an Artin algebra and the pair $(\mathcal{X},\mathcal{Y})$ be a torsion pair in ${\rm{{mod}}}(\Gamma)$. Then the following conditions are equivalent.
  • $(a)$
    $\mathcal{X}$ is functorially finite;
  • $(b)$
    $\mathcal{Y}$ is functorially finite;
  • $(c)$
    $\mathcal{X}=\mathrm{fac}(X)$ for some $X\in{\rm{{mod}}}(\Gamma)$;
  • $(d)$
    $\mathcal{Y}=\mathrm{sub}(Y)$ for some $Y\in{\rm{{mod}}}(\Gamma)$;
  • $(e)$
    $\mathcal{P}(\mathcal{X})$ is a tilting $(\Gamma/\mathrm{Ann}_{\Gamma}(\mathcal{X}))$-module;
  • $(f)$
    $\mathcal{I}(\mathcal{Y})$ is a cotilting $(\Gamma/\mathrm{Ann}_{\Gamma}(\mathcal{Y}))$-module;
  • $(g)$
    $\mathcal{X}=\mathrm{fac}(\mathcal{P}(\mathcal{X}))$;
  • $(h)$
    $\mathcal{Y}=\mathrm{sub}(\mathcal{I}(\mathcal{Y}))$.
1.6. Auslander-Reiten Theory
In this subsection we recall several results in Auslander-Reiten theory.
Let $\Gamma$ be an Artin algebra. The functor
()*$(-)^{*}:={\rm{Hom}}_{\Gamma}(-,\Gamma):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{{\rm{op}}})$
defined by
$M^{*}:={\rm{Hom}}_{\Gamma}(M,\Gamma)$
induces a duality ()*$(-)^{*}:{\rm{add}}(\Gamma)\longrightarrow{\rm{add}}(\Gamma^{{\rm{op}}})$, see e.g. [ARS, Proposition II.4.3].
For a $\Gamma$-module $X$, take a projective presentation $P_{1}\stackrel{{\scriptstyle g}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle f}}{{\longrightarrow}}X\longrightarrow 0.$ Define $\Gamma^{{\rm{op}}}$-module $\mathrm{Tr}X$ by an exact sequence $P_{0}^{*}\stackrel{{\scriptstyle g^{*}}}{{\longrightarrow}}P_{1}^{*}\longrightarrow\mathrm{Tr}X\longrightarrow 0.$ In fact $\mathrm{Tr}X$ is defined as the cokernel of $g^{*}$.
On the other hand, the canonical duality $\mathbf{D}:{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$ induces a duality $\mathbf{D}:\underline{{\rm{{mod}}}}(\Gamma)\longrightarrow\overline{{\rm{{mod}}}}(\Gamma^{\rm{\rm{op}}})$ and we have mutually quasi-inverse equivalences
$\tau:=\mathbf{D}\mathrm{Tr}:\underline{{\rm{{mod}}}}(\Gamma)\stackrel{{\scriptstyle\mathrm{Tr}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma^{{\rm{op}}})\stackrel{{\scriptstyle\mathbf{D}}}{{\longrightarrow}}\overline{{\rm{{mod}}}}(\Gamma)$
$\tau^{-}:=\mathrm{Tr}\mathbf{D}:\overline{{\rm{{mod}}}}(\Gamma)\stackrel{{\scriptstyle\mathbf{D}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma^{{\rm{op}}})\stackrel{{\scriptstyle\mathrm{Tr}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma)$
called the Auslander-Reiten translations. Especially, $\tau$ gives a bijection from isoclasses of indecomposable non-projective $\Gamma$-modules to isoclasses of indecomposable non-injective $\Gamma$-modules, see e.g. [ARS, Propositions IV.1.9 and IV.1.10].

### v2

@@ Abstract.
Let $n$ be a positive integer. In this paper, by using torsion theoretic methods, we show that $n$-Auslander algebras can be characterized by the abelianness of the category of modules with projective dimension less than $n$ and two additional properties, extending the classical Auslander-Tachikawa theorem. By Auslander-Iyama correspondence a categorical characterization of the class of Artin algebras having $n$-cluster tilting modules is obtained.
Since higher Auslander algebras are a special case of higher Auslander-Gorenstein algebras, the results are given in the general setting as extending previous results of Kong.
@@ 1. Introduction
In 1971, Auslander [3] proved a remarkable classical result, called Auslander correspondence, about algebras of finite representation type which introduced a completely new insight to the representation theory of Artin algebras. He showed that there is a bijective correspondence between the set of Morita equivalence classes of Artin algebras $\Lambda$ of finite representation type and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an Auslander algebra; i.e. Artin algebras $\Gamma$ whose global dimensions (denoted by ${\rm{gldim}}(\Gamma)$) are at most two and whose dominant dimensions (denoted by ${\rm{domdim}}(\Gamma)$) are at least two. See [3, Page 52] or [5, Theorem VI.5.7].
For an Artin algebra $\Gamma$, let ${\rm{Mod}}(\Gamma)$ be the category of all left $\Gamma$-modules and ${\rm{{mod}}}(\Gamma)$ be the category of all finitely generated left $\Gamma$-modules. If $\Gamma$ is an Auslander algebra, then the full subcategory $\mathrm{proj}(\Gamma)$ of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated projective $\Gamma$-modules is equivalent to a module category and so is itself an abelian category [3, Page 52].
This gave Tachikawa the motivation to study rings $\Gamma$ such that the full subcategory ${\mathrm{Proj}}(\Gamma)$ of ${\rm{Mod}}(\Gamma)$ consisting of all projective $\Gamma$-modules is abelian. He obtained some characterizations of such rings and proved that these algebras are precisely Auslander algebras [44, Theorem 1]. For Artin algebras, by combining Auslander’s result and Tachikawa’s result, we have the following characterization of Auslander algebras.
@@
In 2007, Iyama in a series of important papers [26, 27] gave a generalization of Auslander correspondence, called higher Auslander correspondence or Auslander-Iyama correspondence. He showed that for every positive integer $n$, there is a bijective correspondence between the set of equivalence classes of $n$-cluster tilting subcategories with additive generators (Definition 6.4) of Artin algebras and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that ${\rm{gldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$, called $n$-Auslander algebras [27, Theorem 0.2].
A natural question is whether there exists a characterization for higher Auslander algebras analogous to the Auslander-Tachikawa Theorem. The first goal of this paper is to answer this question and we have the following characterizations for higher Auslander algebras; see Theorem 6.1.
@@
Here, for an Artin algebra $\Gamma$, we denote by $\Gamma^{{\rm{op}}}$ the opposite algebra of $\Gamma$, by ${\mathrm{Inj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all injective $\Gamma$-modules, by $\mathrm{inj}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated injective $\Gamma$-modules; for every non-negative integer $m$, by ${\mathrm{Proj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all $\Gamma$-modules of projective dimension at most $m$, and by $\mathrm{proj}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules of projective dimension at most $m$. For $m\geq 1$, we write
$^{\perp_{m}}\Gamma:={}^{\perp_{[1,m]}}\Gamma=\{M\in{\rm{{mod}}}(\Gamma)\ |\ {\rm{Ext}}^{i}_{\Gamma}(M,\Gamma)=0,\ \forall\ 1\leq i\leq m\},$
@@
$^{\underline{\perp}_{m}}\Gamma:={}^{\underline{\perp}_{[1,m]}}\Gamma=\{M\in{\rm{Mod}}(\Gamma)\ |\ {\rm{Ext}}^{i}_{\Gamma}(M,\Gamma)=0,\ \forall\ 1\leq i\leq m\};$
@@
Note that only the abelianness property of the category of modules with projective dimension less than $n$ can not characterize $n$-Auslander algebras. For $n\geq 2$, the other conditions are necessary; see Remarks 6.2.
In 1993, Auslander and Solberg [10] established a Gorenstein analogue of Auslander correspondence by using the technique of relative homology developed in [7, 8, 9]. They introduced $\tau$-selfinjective algebras, where $\tau$ is the Auslander-Reiten translation, as a replacement of representation-finite algebras. The corresponding analogue of Auslander algebras are characterized as Artin algebras whose injective dimensions are at most $2$ and whose dominant dimensions are at least $2$, that is, Artin algebras $\Gamma$ such that ${\rm{id}}({}_{\Gamma}\Gamma)\leq 2\leq{\rm{domdim}}(\Gamma)$. Kong named this class of algebras as quasi-Auslander algebras [33]. In his work [33] Kong also proved an extension of Auslander-Tachikawa theorem by characterizing quasi-Auslander algebras as Artin algebras over which the category of finitely generated Gorenstein projective modules is an abelian category.
To obtain a higher analogue of Auslander-Solberg correspondence, Iyama and Solberg [31] introduced pre-cluster tilting subcategories in module categories of Artin algebras which are called $\tau_{n}$-selfinjective algebras as a higher version of $\tau$-selfjinjective algebras. They also gave the notion of $n$-minimal Auslander-Gorenstein algebras as a higher version of quasi-Auslander algebras, say Artin algebras $\Gamma$ with ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq{\rm{domdim}}(\Gamma)$. They showed that there exists a bijective correspondence between Morita equivalence classes of $n$-minimal Auslander-Gorenstein algebras and equivalence classes of finite $n$-precluster tilting subcategories of Artin algebras.
In view of the above Iyama-Solberg correspondence, we could establish a Gorenstein version of Theorem 2 which are also higher versions of Kong’s result [33, Theorem 2.1].
@@
Here for every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$, by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
While the results of Section 3 mainly make use of torsion theoretic methods, in Section 5, we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [11]. Especially, by Corollary 5.4, if $n$ is a positive integer and $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, then we have that
@@ 2.1. Conventions
Let $\Gamma$ be an Artin algebra defined over $R$. We denote by $\mathbf{D}$ the canonical duality ${\rm{Hom}}_{R}(-,\mathbf{E}$ $(R/J_{R})):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$, where $\mathbf{E}(R/J_{R})$ is the injective envelope of $R/J_{R}$ as an $R$-module.
@@
For a module $M\in{\rm{{mod}}}(\Gamma)$ and $i\geq 0$, write $\mathrm{sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{add}(M))$, $\mathrm{fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{add}(M))$, $\mathrm{Sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{Add}(M))$, $\mathrm{Fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{Add}(M))$. In particular, $\mathrm{sub}(M):=\mathrm{sub}^{1}(M)$, $\mathrm{fac}(M)$ $\mathrm{fac}_{1}(M)$, $\mathrm{Sub}(M):=\mathrm{Sub}^{1}(M)$, $\mathrm{Fac}(M):=\mathrm{Fac}_{1}(M)$.
@@ 2.2. Dominant Dimension
The left dominant dimension of the ring $\Gamma$ is defined as the dominant dimension of the left regular module $_{\Gamma}\Gamma$}. Note that the dominant dimension of a right module and the right dominant dimension of a ring are defined similarly. It is well-known that ${\rm{domdim}}({{}_{\Gamma}\Gamma})={\rm{domdim}}(\Gamma_{\Gamma})$ ; see [39, Theorem 4]. So for the rest of the paper, we will denote both left and right dominant dimension of $\Gamma$, by ${\rm{domdim}}(\Gamma)$ and call it the dominant dimension of $\Gamma$.
@@
The inclusion ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$ can be proved exactly in the same way. ∎
@@ 2.3. Torsion Pairs
Let $\Gamma$ be an Artin algebra and $\mathcal{A}:={\rm{Mod}}(\Gamma)$. If $X,Y\in\mathcal{A}$, then we denote the set of morphisms from $X$ to $Y$ in $\mathcal{A}$ by $\mathcal{A}(X,Y)$. A torsion pair in $\mathcal{A}$ is a pair $(\mathcal{X},\mathcal{Y})$ of full subcategories of $\mathcal{A}$ such that $\mathcal{A}(X,Y)=0$ for all $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$, and these two classes are maximal for this property [16, Sections 1 and 2], that is,
@@
A subcategory of $\mathcal{A}$ is the torsion class (resp. the torsion-free class) of some torsion pair if and only if it is closed under quotients, direct sums and extensions (resp. subobjects, direct products and extensions) [16, Theorem 2.3].
A torsion pair $(\mathcal{X},\mathcal{Y})$ in $\mathcal{A}$ is called hereditary if $\mathcal{X}$ is also hereditary, i.e. $\mathcal{X}$ is closed under subobjects (which is equivalent to $\mathcal{Y}$ being closed under taking injective envelopes [16, Theorem 2.9]). For more details about torsion theories in ${\rm{Mod}}(\Gamma)$; see [43, Chapter VI].
We would like to mention that a subcategory of ${\rm{{mod}}}(\Gamma)$ is the torsion class (resp. the torsion-free class) of some torsion pair if and only if it is closed under quotients and extensions (resp. submodules and extensions). For more details about torsion theories in ${\rm{{mod}}}(\Gamma)$; see [1, Subsection 1.1] or [25, Subsection 1.2].
@@
We recall some definitions and notations about (co)tilting modules; see [22, 41].
@@
Let $\Gamma$ be an Artin algebra. A module $T\in{\rm{{mod}}}(\Gamma)$ is said to be a tilting module, if the following three properties are satisfied.
@@
    There is an exact sequence $0\longrightarrow\Gamma\longrightarrow T^{\prime}\longrightarrow T^{\prime\prime}\longrightarrow 0$ with $T^{\prime}$ and $T^{\prime\prime}$ in ${\rm{add}}(T)$.
Dually, a module $T\in{\rm{{mod}}}(\Gamma)$ will be called a cotilting module if $\mathbf{D}(T)$ is a tilting $\Gamma^{\rm op}$-module.